Operations Report – January 31st

Operational Report
Crew 188: Zac Trolley

There is a request by the Hab Director for an electrician to investigate a malfunctioning power plug at her residence.
The generator shut it’s self off at 4am, the assumed fault was the 12 Volt Battery running low. Luckily the crew was up for the lunar eclipse, and were able to swap the batteries out. The second battery lasted less than 12 hours.
A repair crew came and replaces the alternator on the generator. The oil, fuel, and air filters were also changed. The battery is charging at 14V, so I have high confidence that the issue is solved.

Evac plan
• Generator (hours run): 6.1h
• Generator turned off: 10:33 @ SOC 100%
• Generator turned on: 14:00 @SOC 100%

Fuel Levels:
• Diesel: 79%
• Propane: 62%
• Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water Levels:
• Water (trailer): 0 Gallons
• Water (static): 350 Gallons
• Trailer to Static Pump used: No
• Water (loft) – Static to Loft Pump used: Yes
• Meter @ 20:01 130581.4
• Water Used today: 30.5 Gal
• Toilet tank emptied: No

• ATVs were not used today.
• ATVs Used: None
• Oil Added: None
• ATV Fuel Used: None

The Hab Director mentioned that the Rovers required an extension cord that is shorter than 25 feet. We estimate the cords being used now are 50 foot cables. I was able to find a 25 foot extension cable, meaning we could charge one rover at time based on the above specifications. Can we have the appropriate extension cables sent to the Hab in order to charge the Rovers?

Deimos rover used: No
• Hours: 106.9h (Unknown, Director used Rover today, do not have updated numbers)

Spirit rover used: No
• Hours: 17.5h

Opportunity rover used: No
• Hours: 13.5h

Curiosity rover used: No
• Hours: 9.3h

HabCar: Not used

Summary of internet: We purchased the 75 Gig internet token for data in order to support the mission. It is unclear to the crew how the internet tokens operate, and we will monitor the system to ensure we are using is correctly.
Summary of suits and radios: All nominal
Summary of Hab operations: All nominal
Summary of EVA operations: We are still brainstorming ways to combat the helmet fog.
Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.
Summary of Science Dome operations: The science dome is awaiting a portable heater as per the Hab Director.
Summary of RAM operations: Not Operational.
Summary of health and safety issues: There was a report of a brief propane smell today, still awaiting the propane detector to rule out off gassing in the Hab.
Questions, concerns and requests to Mission Support: All nominal

Copyright © The Mars Society. All rights reserved. | Main Site